OpenStudy (anonymous):

sin(sin-1π) = ? i do not remember our lesson on this from friday.

OpenStudy (anonymous):

first off, you don't need to bother with the one in front of pi, and \[\sin(a+b)=\sin(a)\cos(b)+\sin(b)\cos(a)\] keep that one handy. so use this expansion

OpenStudy (anonymous):

sorry that isn't a one, thats inverse sin

OpenStudy (anonymous):

oh well when you have \[sinx = a \] how do you get the value of x? you take the inverse, so that means \[\arcsin(sinx) =sina\] \[acrsin(sinx)=x = sina\]

OpenStudy (anonymous):

so it = π?

OpenStudy (anonymous):

it is not in -π/2 to π/2 though

OpenStudy (anonymous):

yes it is :) from -pi/2 to pi/2 is a full rotation, which is going to include pi

OpenStudy (anonymous):